OpenStudy (anonymous):

Barbara draws pens randomly from a box containing 5 pens of the same shape and size. There is 1 green pen, 3 red pens, and 1 blue pen. She draws 1 red pen and then another red pen without replacing the first one. Find the probability of drawing 1 red pen followed by another red pen, and show the equation used.

OpenStudy (anonymous):

1/5 + 1/4...or am I doing it incorrectly? It is multiplying isnt it?

OpenStudy (anonymous):

@Valpey

OpenStudy (valpey):

It is multiplying, but you have the right idea although you want 3 red pens in the numerator to start, not 1.

OpenStudy (valpey):

\[\frac{red\ pens}{all\ pens} *\frac{remaining\ red\ pens}{all\ remainin\ pens} = \frac{3}{5}*\frac{2}{4} = \]

OpenStudy (anonymous):

OKAY! Thanks so much for your help! :)